\begin{thebibliography}{99}

%\cite{:2011zb}
\bibitem{:2011zb}
  T.~Adam {\it et al.} [ OPERA Collaboration ],
  ``Measurement of the neutrino velocity with the OPERA detector in the CNGS beam,''
    [arXiv:1109.4897 [hep-ex]].


%\cite{Bernabei:2008yi}
\bibitem{Bernabei:2008yi}
  R.~Bernabei {\it et al.} [ DAMA Collaboration ],
  ``First results from DAMA/LIBRA and the combined results with DAMA/NaI,''
  Eur.\ Phys.\ J.\  {\bf C56}, 333-355 (2008).
  [arXiv:0804.2741 [astro-ph]].
%\cite{Cacciapaglia:2011ax}
\bibitem{Cacciapaglia:2011ax}
  G.~Cacciapaglia, A.~Deandrea, L.~Panizzi,
  ``Superluminal neutrinos in long baseline experiments and SN1987a,''
  [arXiv:1109.4980 [hep-ph]].

%\cite{AmelinoCamelia:2011dx}
\bibitem{AmelinoCamelia:2011dx}
  G.~Amelino-Camelia, G.~Gubitosi, N.~Loret, F.~Mercati, G.~Rosati, P.~Lipari,
  ``OPERA-reassessing data on the energy dependence of the speed of neutrinos,''
  [arXiv:1109.5172 [hep-ph]].

%\cite{Dvali:2011mn}
\bibitem{Dvali:2011mn}
  G.~Dvali, A.~Vikman,
  ``Price for Environmental Neutrino-Superluminality,''
  [arXiv:1109.5685 [hep-ph]].

%\cite{Ciborowski:2011ka}
\bibitem{Ciborowski:2011ka}
  J.~Ciborowski, J.~Rembielinski,
  ``Comments on the recent velocity measurement of the muon neutrinos by the OPERA Collaboration,''
  [arXiv:1109.5599 [hep-ex]].

%\cite{Klinkhamer:2011mf}
\bibitem{Klinkhamer:2011mf}
  F.~R.~Klinkhamer,
  ``Superluminal muon-neutrino velocity from a Fermi-point-splitting model of Lorentz violation,''
  [arXiv:1109.5671 [hep-ph]].

%\cite{Autiero:2011hh}
\bibitem{Autiero:2011hh}
  D.~Autiero, P.~Migliozzi, A.~Russo,
  ``The neutrino velocity anomaly as an explanation of the missing observation of neutrinos in coincidence with GRB,''
  [arXiv:1109.5378 [hep-ph]].

%\cite{Alexandre:2011kr}
\bibitem{Alexandre:2011kr}
  J.~Alexandre,
  ``Lifshitz-type Quantum Field Theories in Particle Physics,''
  [arXiv:1109.5629 [hep-ph]].

%\cite{Giudice:2011mm}
\bibitem{Giudice:2011mm}
  G.~F.~Giudice, S.~Sibiryakov, A.~Strumia,
  ``Interpreting OPERA results on superluminal neutrino,''
  [arXiv:1109.5682 [hep-ph]].

%\cite{Gubser:2011mp}
\bibitem{Gubser:2011mp}
  S.~S.~Gubser,
  ``Superluminal neutrinos and extra dimensions: constraints from the null energy condition,''
  [arXiv:1109.5687 [hep-th]].

%\cite{Svozil:2011mu}
\bibitem{Svozil:2011mu}
  K.~Svozil,
  ``Neutrino dispersion relation changes due to radiative corrections as the origin of faster-than-light-in-vacuum propagation in a medium,''
  [arXiv:1109.5411 [physics.gen-ph]].

%\cite{Cahill:2011mt}
\bibitem{Cahill:2011mt}
  K.~Cahill,
  ``Fast Light, Fast Neutrinos?,''
  [arXiv:1109.5357 [physics.gen-ph]].

%\cite{Cardone:2011ms}
\bibitem{Cardone:2011ms}
  F.~Cardone, R.~Mignani, A.~Petrucci,
  ``Neutrinos superluminality and Local Lorentz Invariance,''
  [arXiv:1109.5289 [physics.gen-ph]].

%\cite{Mann:2011rd}
\bibitem{Mann:2011rd}
  R.~B.~Mann, U.~Sarkar,
  ``Superluminal neutrinos at the OPERA?,''
  [arXiv:1109.5749 [hep-ph]].

%\cite{Alicki:2011qw}
\bibitem{Alicki:2011qw}
  R.~Alicki,
  ``A possible statistical mechanism of anomalous neutrino velocity in OPERA experiment?,''
  [arXiv:1109.5727 [physics.data-an]].

%\cite{Pfeifer:2011ve}
\bibitem{Pfeifer:2011ve}
  C.~Pfeifer, M.~N.~R.~Wohlfarth,
  ``Beyond the speed of light on Finsler spacetimes,''
  [arXiv:1109.6005 [gr-qc]].

%\cite{Li:2011ue}
\bibitem{Li:2011ue}
  M.~Li, T.~Wang,
  ``Mass-dependent Lorentz Violation and Neutrino Velocity,''
  [arXiv:1109.5924 [hep-ph]].

%\cite{Drago:2011ua}
\bibitem{Drago:2011ua}
  A.~Drago, I.~Masina, G.~Pagliara, R.~Tripiccione,
  ``The Hypothesis of Superluminal Neutrinos: comparing OPERA with other Data,''
  [arXiv:1109.5917 [hep-ph]].

%\cite{Contaldi:2011zm}
\bibitem{Contaldi:2011zm}
  C.~R.~Contaldi,
  ``The OPERA neutrino velocity result and the synchronisation of clocks,''
  [arXiv:1109.6160 [hep-ph]].

%\cite{Oikonomou:2011zp}
\bibitem{Oikonomou:2011zp}
  V.~K.~Oikonomou,
  ``The 2d Gross-Neveu Model for Pseudovector Fermions and Tachyonic Mass Generation,''
  [arXiv:1109.6170 [hep-th]].

%\cite{Konoplya:2011ag}
\bibitem{Konoplya:2011ag}
  R.~A.~Konoplya,
  ``Superluminal neutrinos and the tachyon's stability in the rotating Universe,''
  [arXiv:1109.6215 [hep-th]].

%\cite{Iorio:2011ay}
\bibitem{Iorio:2011ay}
  L.~Iorio,
  ``Environmental fifth-force hypothesis for the OPERA superluminal neutrino phenomenology: constraints from orbital motions around the Earth,''
  [arXiv:1109.6249 [gr-qc]].

%\cite{Anacleto:2011bv}
\bibitem{Anacleto:2011bv}
  M.~A.~Anacleto, F.~A.~Brito, E.~Passos,
  ``Supersonic Velocities in Noncommutative Acoustic Black Holes,''
  [arXiv:1109.6298 [hep-th]].

%\cite{Kehagias:2011cb}
\bibitem{Kehagias:2011cb}
  A.~Kehagias,
  ``Relativistic Superluminal Neutrinos,''
  [arXiv:1109.6312 [hep-ph]].

%\cite{Alexandre:2011bu}
\bibitem{Alexandre:2011bu}
  J.~Alexandre, J.~Ellis, N.~E.~Mavromatos,
  ``On the Possibility of Superluminal Neutrino Propagation,''
  [arXiv:1109.6296 [hep-ph]].

%\cite{Magueijo:2011xy}
\bibitem{Magueijo:2011xy}
  J.~Magueijo,
  ``Neutrino oscillations and superluminal propagation,''
  [arXiv:1109.6055 [hep-ph]].

%\cite{Hannestad:2011bj}
\bibitem{Hannestad:2011bj}
  S.~Hannestad, M.~S.~Sloth,
  ``Apparent faster than light propagation from light sterile neutrinos,''
  [arXiv:1109.6282 [hep-ph]].

%\cite{GonzalezMestres:2011cd}
\bibitem{GonzalezMestres:2011cd}
  L.~Gonzalez-Mestres,
  ``Comments on the recent result of the 'Measurement of the neutrino velocity with the OPERA detector in the CNGS beam',''
  [arXiv:1109.6308 [physics.gen-ph]].

%\cite{Pankovic:2011cc}
\bibitem{Pankovic:2011cc}
  V.~Pankovic,
  ``Is OPERA Neutrino Superluminal Propagation similar to Gain-Assisted Superluminal Light Propagation,''
  [arXiv:1109.6121 [physics.gen-ph]].

%\cite{Nicolaidis:2011eq}
\bibitem{Nicolaidis:2011eq}
  A.~Nicolaidis,
  ``Neutrino Shortcuts in Spacetime,''
  [arXiv:1109.6354 [hep-ph]].

%\cite{Cohen:2011hx}
\bibitem{Cohen:2011hx}
  A.~G.~Cohen, S.~L.~Glashow,
  ``New Constraints on Neutrino Velocities,''
  [arXiv:1109.6562 [hep-ph]].

%\cite{Klinkhamer:2011iz}
\bibitem{Klinkhamer:2011iz}
  F.~R.~Klinkhamer, G.~E.~Volovik,
  ``Superluminal neutrino and spontaneous breaking of Lorentz invariance,''
  [arXiv:1109.6624 [hep-ph]].

%\cite{GonzalezMestres:2011jc}
\bibitem{GonzalezMestres:2011jc}
  L.~Gonzalez-Mestres,
  ``Astrophysical consequences of the OPERA superluminal neutrino,''
  [arXiv:1109.6630 [physics.gen-ph]].

%\cite{Ciuffoli:2011ji}
\bibitem{Ciuffoli:2011ji}
  E.~Ciuffoli, J.~Evslin, J.~Liu, X.~Zhang,
  ``OPERA and a Neutrino Dark Energy Model,''
  [arXiv:1109.6641 [hep-ph]].

%\cite{Bi:2011nd}
\bibitem{Bi:2011nd}
  X.~-J.~Bi, P.~-F.~Yin, Z.~-H.~Yu, Q.~Yuan,
  ``Constraints and tests of the OPERA superluminal neutrinos,''
  [arXiv:1109.6667 [hep-ph]].

%\cite{Wang:2011sz}
\bibitem{Wang:2011sz}
  P.~Wang, H.~Wu, H.~Yang,
  ``Superluminal neutrinos and domain walls,''
  [arXiv:1109.6930 [hep-ph]].

%\cite{Henri:2011wu}
\bibitem{Henri:2011wu}
  G.~Henri,
  ``A simple explanation of OPERA results without strange physics,''
  [arXiv:1110.0239 [hep-ph]].

%\cite{Cowsik:2011wv}
\bibitem{Cowsik:2011wv}
  R.~Cowsik, S.~Nussinov, U.~Sarkar,
  ``Superluminal Neutrinos at OPERA Confront Pion Decay Kinematics,''
  [arXiv:1110.0241 [hep-ph]].

%\cite{Torrealba:2011ww}
\bibitem{Torrealba:2011ww}
  R.~Torrealba,
  ``Using an Einstein's idea to explain OPERA faster than light neutrinos,''
  [arXiv:1110.0243 [hep-ph]].

%\cite{Dass:2011yj}
\bibitem{Dass:2011yj}
  N.~D.~H.~Dass,
  ``OPERA, SN1987a and energy dependence of superluminal neutrino velocity,''
  [arXiv:1110.0351 [hep-ph]].

%\cite{Winter:2011zf}
\bibitem{Winter:2011zf}
  W.~Winter,
  ``How large is the fraction of superluminal neutrinos at OPERA?,''
  [arXiv:1110.0424 [hep-ph]].

%\cite{Anber:2011vk}
\bibitem{Anber:2011vk}
  M.~M.~Anber, J.~F.~Donoghue,
  ``Limiting velocities as running parameters and superluminal neutrinos,''
  [arXiv:1110.0132 [hep-ph]].

%\cite{Franklin:2011ws}
\bibitem{Franklin:2011ws}
  J.~Franklin,
  ``Superluminal neutrinos,''
  [arXiv:1110.0234 [physics.gen-ph]].

%\cite{Li:2011xp}
\bibitem{Li:2011xp}
  S.~-Y.~Li,
  ``OPERA Collaboration have observed phase speed of neutrino wave function,''
  [arXiv:1110.0302 [hep-ph]].

%\cite{Carmona:2011zg}
\bibitem{Carmona:2011zg}
  J.~M.~Carmona, J.~L.~Cortes,
  ``Constraints from Neutrino Decay on Superluminal Velocities,''
  [arXiv:1110.0430 [hep-ph]].

%\cite{Li:2011zm}
\bibitem{Li:2011zm}
  T.~Li, D.~V.~Nanopoulos,
  ``Background Dependent Lorentz Violation from String Theory,''
  [arXiv:1110.0451 [hep-ph]].

%\cite{Aref'eva:2011zp}
\bibitem{Aref'eva:2011zp}
  I.~Y.~.Aref'eva, I.~V.~Volovich,
  ``Superluminal Dark Neutrinos,''
  [arXiv:1110.0456 [hep-ph]].

%\cite{Lust:2011fx}
\bibitem{Lust:2011fx}
  D.~Lust, M.~Petropoulos,
  ``Comment on superluminality in general relativity,''
  [arXiv:1110.0813 [gr-qc]].

%\cite{Broda:2011du}
\bibitem{Broda:2011du}
  B.~Broda,
  ``A classical model explaining the OPERA velocity paradox,''
  [arXiv:1110.0644 [physics.data-an]].

%\cite{AmelinoCamelia:2011bz}
\bibitem{AmelinoCamelia:2011bz}
  G.~Amelino-Camelia, L.~Freidel, J.~Kowalski-Glikman, L.~Smolin,
  ``OPERA neutrinos and relativity,''
  [arXiv:1110.0521 [hep-ph]].

%\cite{Brustein:2011fh}
\bibitem{Brustein:2011fh}
  R.~Brustein, D.~Semikoz,
  ``Apparent superluminal neutrino propagation caused by nonlinear coherent interactions in matter,''
  [arXiv:1110.0762 [hep-ph]].

%\cite{Maccione:2011fr}
\bibitem{Maccione:2011fr}
  L.~Maccione, S.~Liberati, D.~M.~Mattingly,
  ``Violations of Lorentz invariance in the neutrino sector after OPERA,''
  [arXiv:1110.0783 [hep-ph]].

%\cite{Davoudiasl:2011gb}
\bibitem{Davoudiasl:2011gb}
  H.~Davoudiasl, T.~G.~Rizzo,
  ``Testing the OPERA Superluminal Neutrino Anomaly at the LHC,''
  [arXiv:1110.0821 [hep-ph]].

%\cite{Goldberg:2011mc}
\bibitem{Goldberg:2011mc}
  J.~Goldberg,
  ``Comment on Jerrold Franklin, Superluminal neutrinos, arXiv:1110.0304,''
  [arXiv:1110.0970 [hep-ex]].

%\cite{Naumov:2011mm}
\bibitem{Naumov:2011mm}
  D.~V.~Naumov, V.~A.~Naumov,
  ``Neutrino Velocity Anomalies: A Resolution without a Revolution,''
  [arXiv:1110.0989 [hep-ph]].

%\cite{Oda:2011kh}
\bibitem{Oda:2011kh}
  I.~Oda, H.~Taira,
  ``Superluminal Neutrinos from Gauge Field,''
  [arXiv:1110.0931 [hep-ph]].

%\cite{Kraniotis:2011sr}
\bibitem{Kraniotis:2011sr}
  G.~V.~Kraniotis,
  ``Exact deflection of a Neutral-Tachyon in the Kerr's Gravitational field,''
  [arXiv:1110.1223 [gr-qc]].

%\cite{Sanchez:2011tu}
\bibitem{Sanchez:2011tu}
  P.~A.~Sanchez, M.~Anabitarte, M.~Bellini,
  ``Dirac equation for massive neutrinos in a non-rotating SdS black hole from a 5D vacuum,''
  [arXiv:1110.1300 [hep-th]].

%\cite{Ahluwalia:2011rg}
\bibitem{Ahluwalia:2011rg}
  D.~V.~Ahluwalia, S.~P.~Horvath, D.~Schritt,
  ``Probing neutrino masses with neutrino-speed experiments,''
  [arXiv:1110.1162 [hep-ph]].

%\cite{Mecozzi:2011tc}
\bibitem{Mecozzi:2011tc}
  A.~Mecozzi, M.~Bellini,
  ``Superluminal group velocity of neutrinos,''
  [arXiv:1110.1253 [hep-ph]].

%\cite{Xue:2011tz}
\bibitem{Xue:2011tz}
  S.~-S.~Xue,
  ``Do high-energy neutrinos travel faster than photons in a discrete space-time?,''
  [arXiv:1110.1317 [hep-ph]].

%\cite{Moffat:2011ue}
\bibitem{Moffat:2011ue}
  J.~W.~Moffat,
  ``Bimetric Relativity and the Opera Neutrino Experiment,''
  [arXiv:1110.1330 [hep-ph]].

\bibitem{a:2011zb}
  T.~Adam {\it et al.} [ OPERA Collaboration ],
  ``Measurement of the neutrino velocity with the OPERA detector in the CNGS beam,''
    [arXiv:1109.4897 [hep-ex]].

%\cite{Morris:1988tu}
\bibitem{Morris:1988tu}
  M.~S.~Morris, K.~S.~Thorne, U.~Yurtsever,
  ``Wormholes, Time Machines, and the Weak Energy Condition,''
  Phys.\ Rev.\ Lett.\  {\bf 61}, 1446-1449 (1988).
  
%\cite{Kalbfleisch:1979rm}
\bibitem{Kalbfleisch:1979rm}
  G.~R.~Kalbfleisch, N.~Baggett, E.~C.~Fowler, J.~Alspector,
  ``Experimental Comparison Of Neutrino, Anti-neutrino, And Muon Velocities,''
  Phys.\ Rev.\ Lett.\  {\bf 43}, 1361 (1979).

%\cite{Blum:2011jf}
\bibitem{Blum:2011jf}
  K.~Blum,
  ``DAMA vs. the annually modulated muon background,''
    [arXiv:1110.0857 [astro-ph.HE]].

\bibitem{wip}
G.X.\ Gilbert-Thorple, J.\ Jesus, in preparation
  

\end{thebibliography}

\documentclass{scrartcl}
\usepackage{amsmath}
\usepackage{graphicx}

\title{Common origin of superluminal neutrinos and DAMA annual modulation}
\author{ Gerald Xavier Gilbert-Thorple$^*$, J\^os\'e J.\ Jesus$^\dagger$\\ 
  \textit{$^*$Kavli University of Southern North Dakota,
Hoople, ND, USA}\\
  \textit{$^\dagger$Beijing University, Athens Campus, 
Athens, Greece, PRC}
}
\date{}

\begin{document}
\maketitle

\begin{abstract}
We suggest that the recently observed superluminal neutrino
propagation observed by OPERA may have an environmental origin
that can simultaneously explain the DAMA/LIBRA annual modulation.
Our proposal can be unambiguously tested by the prediction that
OPERA will also observe an annual modulation in the speed of
neutrinos.
Experiments geographically distant from Gran Sasso will however not be
able to observe these effects.  We do not rely upon any violation of
Lorentz symmetry except for spontaneous breaking of 
translational invariance.  We discuss implications for the CRESST
experiment, also located at Gran Sasso.

\end{abstract}

In a {\it tour de maitre} of striking disproportions, the physics world
and indeed fragments of the general public have been astonished by the 
long-awaited overthrow of relativistic despotism implied by the OPERA
experiment's recent detection of muon neutrinos traveling faster than
light between CERN and Gran Sasso National Laboratory (LNGS)
\cite{:2011zb}.   Before declaring such a victory
however, we should be careful to rule out more mundane explanations
such as the one proposed in the present paper (see below).  It is
intriguing that at the same time (plus or minus ten years) the
DAMA/LIBRA experiment, seemingly coincidentally located at 
Gran Sasso,
has been reporting equally convincing evidence for the annual
modulation of unidentifiable backgrounds in their detector
\cite{Bernabei:2008yi}, which
could include the impotent interactions of dark matter particles blowing
through their apparatus at highly subluminal speeds.  

The dramatic discovery of OPERA has been met with a remarkable paucity
of plausible explanations or criticisms from the theory community 
\cite{AmelinoCamelia:2011dx}-\cite{Moffat:2011ue}.  In a remarkably
cunning bit of sophistry, Cohen {\it et al.} \cite{Cohen:2011hx}
argued that superluminal neutrinos would lose too much energy via
decay to be consistent with the observation, but this hypothesis is
clearly ruled out by the observations of ref.\ \cite{a:2011zb}. 
Similarly ref.\  \cite{Giudice:2011mm} has pointed out theoretical
difficulties for superluminal propagation, but their analysis is not
self-consistent since the authors clearly knew about the experimental
results before they were reported, showing evidence for superluminal
propagation of information.

Our observation is inspired by the curious proposal of ref.\ 
\cite{Dvali:2011mn}, which claimed that an extra graviton polarization
sourced by the earth could account for the superluminal motion. This
claim seems absurd since as is well known, passage through a
gravitational potential can only make particles travel slower, not
faster than light\footnote{We thank Dr.\ G.D.\ Morose for charming
diatribes on this subject}, unless the source has unusual properties.
Here we posit the existence of a source that violates the weak null
energy condition (WNEC),\footnote{We remind the reader that ``null''
in the context of WNEC modifies ``condition'' rather than ``energy''}
which is also a requirement for sourcing traversible wormhole
solutions in general relativity \cite{Morris:1988tu}.  A wormhole
with one end located in close proximity to LNGS would be consistent
with the explanation of ref.\ \cite{Dvali:2011mn}, while removing 
the objections to a power-law behavior for the energy-dependence
of the superluminal effect, since the experiments which would
seemingly contradict this hypothesis \cite{Kalbfleisch:1979rm}
were not located at LNGS.

Emanating from this wormhole will be a number of objects which do not
get very far inside the Gran Sasso Mountain, but in addition we expect a
contribution to the dark matter wind that can have a significant
impact on the annual modulation observed by DAMA.  This observation is
especially timely since it has been recently pointed out that there is
a strong correlation between the DAMA observations and the annually
modulated flux of atmospheric neutrinos \cite{Blum:2011jf}. It is
important to notice that the extra component to the dark matter wind
emanating from the wormhole will also be annually modulated by the
stretching of the wormhole over the course of a year (except in the
unlikely event that the other end of the wormhole lies along the  axis
of rotation of the earth's orbit).  We believe that this effect should
exactly compensate the atmospheric muon background not taken into
account by DAMA's analysis, meaning that their net modulation comes
from the ordinary (non-wormhole) component of the dark matter wind. 
Here we have some connection with the proposal of ref.\ 
\cite{Hannestad:2011bj}, which fantasized the propagation of sterile
 neutrinos mixing with $\nu_\mu$
in warped extra dimensions; our wormhole too could be traversing in
part the extra dimensions.  However unlike ref.\
\cite{Hannestad:2011bj} we do not need to assume that the mixing angle
of $\nu_\mu$ with the sterile neutrino is simultaneously both large and
small.

Unlike other proposed explanations of the OPERA anomaly, our proposal
is fully testable, since the annual distortions of the wormhole will
change the gravitational potential transmitted by the extra graviton
polarization, leading to annual modulations in the $\nu_\mu$ velocity
anomaly of one part in $10^{10}$, which could be measured by OPERA
given a sufficiently long period of data collection and analysis.  In addition,
there are implications for the CRESST experiment which is also located
at LNGS.  These and other details of the present analysis, complete
with equations, will be discussed in a forthcoming publication 
\cite{wip} now that we have laid claim to our idea.

\section*{Acknowledgment}

We thank members of the OPERA and DAMA collaborations for candid
discussions admitting the probability of our explanation.

\end{document}